# Principal twistor models and asymptotic hyperkähler metrics

### 04 Mar 2026 v1

@@ 1. Introduction
A hyperkähler metric is defined as a metric that is Kähler with respect to three complex structures satisfying the quaternionic relations. In this case, we naturally obtain a deformation of holomorphic symplectic manifolds indexed by the Riemann sphere $\mathbb{P}^{1}$, endowed with a real structure and a family of rational curves called twistor lines; this collection of data is referred to as the twistor space. Conversely, when a twistor space is given, the hyperkähler metric can be reconstructed from that information [5]. From this viewpoint, Hitchin [4] constructed ALE gravitational instantons, which are typical examples of algebraic hyperkähler metrics with asymptotic behavior.
@@
We define a twistor model to be a structure satisfying all conditions of a twistor space except for the existence of twistor lines. As a generalization of twistor models, we define a new notion called the principal twistor model, which consists of a deformation of holomorphic symplectic manifolds indexed by a vector bundle $\mathcal{C}(2)$ over $\mathbb{P}^{1}$ and a real structure on it. For any choice of a real section $s$ of the vector bundle $\mathcal{C}(2)$ (cf. Def. 3.8), we obtain a twistor model $Z_{s}$ by slicing the principal twistor model along the section $s$ (cf. Prop. 3.9). In other words, the twistor model $Z_{s}$ is given by restricting the principal twistor model to the real section $s$. As will be explained later, the principal twistor model provides a natural framework for studying the moduli of algebraic hyperkähler structures asymptotic to an algebraic hyperkähler cone (Cor. 1.5). In this framework, the asymptotic behavior of the metric can be interpreted as a degeneration of twistor spaces (cf. Remark 4.11).
In this paper, as the asymptotic hyperkähler cone, we consider a conical symplectic variety $X$ with a good triple (cf. Def. 2.5) consisting of a good $\mathbb{C}^{*}$-action, a holomorphic symplectic form, and a quaternionic structure. As will be explained later, this condition is necessary for a conical symplectic variety to admit an algebraic hyperkähler cone metric (Theorem 1.3). We further assume that the variety $X$ admits a projective crepant resolution $Y$ (which we simply call a crepant resolution).
To study the geometry of $Y$, we use the theory of Poisson deformations. A Poisson deformation is a deformation of a complex manifold equipped with a compatible deformation of the Poisson bracket, which corresponds to a relative holomorphic symplectic form. By a theorem of Namikawa [14], the crepant resolution $Y$ admits the universal Poisson deformation $\mathcal{Y}\to\mathcal{C}$, which captures all such deformations, and its base space $\mathcal{C}$ is naturally isomorphic to $H^{2}(Y;\mathbb{C})$. Based on the idea of Bielawski and Foscolo [2] of gluing this universal Poisson deformation via the $\mathbb{C}^{*}$-action, a principal twistor model with $Y$ as its central fiber is constructed (Prop. 1.1; see Figures 2 and 2 on page 2).
We present three main results of this paper. First, we state that the principal twistor model is naturally constructed from a good triple of a conical symplectic variety. This result is shown by arguments on formal universal Poisson deformations and applying a $\mathbb{C}^{*}$-gluing construction (cf. §3.2):
@@
Let $X$ be a conical symplectic variety with a good triple, and assume that it admits a crepant resolution $Y$. Let $\mathcal{Y}\rightarrow\mathcal{C}$ denote the universal Poisson deformation of $Y$, where the base space $\mathcal{C}$ is isomorphic to $H^{2}(Y;\mathbb{C})$. Then, the good triple naturally extends to the triple on $\mathcal{Y}$. Furthermore, by applying the $\mathbb{C}^{*}$-gluing construction to this triple, the principal twistor model $\mathcal{Y}(1)\rightarrow\mathcal{C}(2)$ is naturally constructed.
@@
The extension of the data other than the quaternionic structure to the universal Poisson deformation is due to Namikawa [13, 16]. In this work, we show that the quaternionic structure is similarly extendable.
The following theorem on the universality of the principal twistor model is the central result of this paper. This theorem states that any twistor space corresponding to an algebraic hyperkähler metric asymptotic to a hyperkähler cone is uniquely obtained by slicing the principal twistor model along a real section. This result is shown by arguments on the universality of the universal Poisson deformation and the conditions of the asymptotic behavior of the metric:
@@
Let $X$ be a conical symplectic variety with a good triple, and assume that there exists a hyperkähler cone metric $g_{0}$ on the regular locus $X_{\mathrm{reg}}$. Then, the metric $g_{0}$ naturally defines the good triple on $X$. Assume further that $X$ admits a crepant resolution $Y$. Let $\mathcal{Y}(1)\rightarrow\mathcal{C}(2)$ be the principal twistor model constructed in Proposition 1.1 for this good triple. Consider an algebraic hyperkähler metric $g$ on $Y$ (cf. Def. 3.29) that is asymptotic at infinity to the metric $g_{0}$ (cf. Def. 3.37). Then, there exists a unique real section $s$ of the vector bundle $\mathcal{C}(2)$ such that the twistor space $Z$ corresponding to the metric $g$ is isomorphic, as a twistor model, to the model $Z_{s}$ obtained by slicing the principal twistor model $\mathcal{Y}(1)$ along the real section $s$.
@@
Finally, as a corollary of Theorem 1.3, we present an application to the moduli space of hyperkähler structures with asymptotic behavior. This result is obtained by combining an analysis of twistor lines with the asymptotic conditions on the metric:
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and that $X$ has a crepant resolution $Y$. Let $\mathcal{M}$ be the moduli space of hyperkähler structures $(Y,g,I,J,K)$ defined as follows:
@@
The organization of the paper is as follows. In §2, we introduce conical symplectic varieties with a good triple and show that the triple extends to the universal Poisson deformation of its crepant resolution. In §3, we define the principal twistor model and explain its construction after introducing the $\mathbb{C}^{*}$-gluing construction. We also prove the universality of the principal twistor model. In §4, as an application of this work, we study the moduli of hyperkähler structures with asymptotic behavior. We also introduce metrics constructed via hyperkähler quotients and QALE hyperkähler metrics as concrete examples to which this work applies under certain assumptions. Finally, we conclude the paper by discussing future directions.
@@ 2.1. Definitions
Next, we define a quaternionic structure on a complex manifold equipped with a $\mathbb{C}^{*}$-action.
@@
An anti-holomorphic automorphism $j$ on a complex manifold $X$ with a $\mathbb{C}^{*}$-action $\lambda$ is called a quaternionic structure on $X$ if it satisfies
@@
    The vector space $\mathbb{C}^{2n}$ admits a standard good triple as follows: For standard coordinates $(x_{1},\ldots,x_{n};y_{1},\ldots,y_{n})$,
@@
    Therefore, $X$ naturally inherits the good triple from the standard good triple on $\mathbb{C}^{2n}$.
@@
2.2. Crepant resolutions
In this subsection, we show that when a conical symplectic variety with a good triple admits a crepant resolution, the good triple naturally lifts to the resolution. That is, we prove the following proposition:
Proposition 2.10.
Let $X$ be a conical symplectic variety with a good triple. Assume that $X$ admits a crepant resolution $Y$. Then, the good triple on $X$ naturally lifts to a triple $(\lambda_{Y},\omega_{Y},j_{Y})$ on the crepant resolution $Y$.
Remark 2.11.
Throughout this paper, a crepant resolution always means a projective crepant resolution.
First, we show the claim for the case of quotient spaces with Kleinian singularities. Consider a vector space $V=\mathbb{C}^{2}$ and a finite group $\Gamma<Sp(1)=SU(2)$, and consider the quotient space $V/\Gamma. XX$ admits a minimal resolution $Y$, which is the unique crepant resolution of $X$ up to isomorphism. The minimal resolution is constructed via the $\Gamma$-Hilbert scheme:
Theorem 2.12 (Ito and Nakamura [6]).
Let $V=\mathbb{C}^{2}$ and $\Gamma<SU(2)$ be a finite group. Consider the quotient space $X=V/\Gamma$ with a Kleinian singularity. Let $n$ be the order of the finite group $\Gamma$. Let $V^{(n)}$ be the $n$-th symmetric product and $V^{[n]}$ the Hilbert scheme of $n$ points. Consider the crepant resolution given by the Hilbert-Chow map
$\pi:V^{[n]}\rightarrow V^{(n)}.$
The action of $\Gamma$ on $V$ naturally extends to actions on $V^{(n)}$ and $V^{[n]}$. Let $(V^{(n)})^{\Gamma}$ and $(V^{[n]})^{\Gamma}$ be the fixed point sets of these actions, respectively. Then, we naturally obtain the restriction of $\pi$:
$\pi^{\Gamma}:(V^{[n]})^{\Gamma}\rightarrow(V^{(n)})^{\Gamma}.$
Let $\mathrm{Hilb}^{\Gamma}(V)$ be the principal connected component of $(V^{[n]})^{\Gamma}$. Then,
$\pi^{\Gamma}:\mathrm{Hilb}^{\Gamma}(V)\rightarrow(V^{(n)})^{\Gamma}$
gives the minimal resolution of $(V^{(n)})^{\Gamma}\simeq X$.
We establish the following lemma.
Lemma 2.13.
Let $V=\mathbb{C}^{2}$, $V^{(n)}$ be the $n$-th symmetric product, and $V^{[n]}$ be the Hilbert scheme of $n$ points. Then, the standard triple on $V$ naturally induces the good triple $(\lambda,\omega,j)$ on the symmetric product $V^{(n)}$. Furthermore, this triple lifts to the Hilbert scheme $V^{[n]}$ via the Hilbert-Chow map
$\pi:V^{[n]}\rightarrow V^{(n)}.$
Proof.
The fact that the holomorphic symplectic $2$-form $\omega$ lifts to $V^{[n]}$ is precisely the result of Beauville and Fujiki [12, Theorem 1.17]. We show that the $\mathbb{C}^{*}$-action and the quaternionic structure lift to $V^{[n]}$. Consider the singular set $V_{1}^{(n)}$ on the symmetric product $V^{(n)}$ defined by $n$ points $\{p_{i}\}$ satisfying $p_{1}=p_{2}$ and $p_{i}\neq p_{1}\ (i\geq 3)$. A neighborhood of a point in $V_{1}^{(n)}$ can be written as the product of an $A_{1}$-singularity and an open set in $\mathbb{C}^{2n-2}$. Let $V^{(n)}_{*}=V^{(n)}_{\mathrm{reg}}\sqcup V_{1}^{(n)}$, and let $V^{[n]}_{*}$ be its preimage under the Hilbert-Chow map $\pi$. This provides a crepant resolution of $V^{(n)}_{*}$. Since the minimal resolution of an $A_{1}$-singularity is given by the blow-up at the origin, the good triple naturally lifts. Thus, the triple lifts to $V^{[n]}_{*}$. Since the complement of $V^{[n]}_{*}$ has codimension at least 2 in $V^{[n]}$, the claim follows from Hartogs’ extension theorem. ∎
Corollary 2.14.
Let $V=\mathbb{C}^{2}$ and $\Gamma<SU(2)$ be a finite group. Consider the quotient space $X=V/\Gamma$ with a Kleinian singularity. Then, the standard good triple $(\lambda,\omega,j)$ on $X$ naturally lifts to the triple on the minimal resolution $Y$ of $X$.
Proof.
Let $n$ be the order of $\Gamma$. By Lemma 2.13, the Hilbert scheme $V^{[n]}$ admits the good triple. Furthermore, similar to Example 2.8 (2), the good triple is compatible with the action of $\Gamma$ on $V^{[n]}$. Therefore, the good triple on $X\simeq(V^{(n)})^{\Gamma}$ naturally lifts to the triple on the minimal resolution $Y\simeq\mathrm{Hilb}^{\Gamma}(V)\subset(V^{[n]})^{\Gamma}$. ∎
With the above preparation, we prove Proposition 2.10.
Proof of Proposition 2.10.
Let $\pi:Y\rightarrow X$ be a crepant resolution. Since $\pi$ is a symplectic resolution, the holomorphic symplectic $2$-form lifts to $Y$. We show that the $\mathbb{C}^{*}$-action $\lambda_{X}$ and the quaternionic structure $j_{X}$ on $X$ lift to $Y$. Recall that by Kaledin [8], the symplectic variety $X$ admits a natural stratification $X=X_{0}\supset X_{1}\supset X_{2}\supset\cdots$ defined inductively by $X_{i+1}=\mathrm{Sing}(X_{i})$, such that each open stratum $X_{i}\setminus X_{i+1}$ is a smooth holomorphic symplectic manifold. (The connected components of these strata correspond to the symplectic leaves of $X$.) Due to the existence of this symplectic stratification, it automatically follows that the codimension of $X_{2}$ in $X$ is at least $4$. Consider the open set $U\coloneqq X\setminus X_{2}$, which is preserved by the actions $\lambda_{X}$ and $j_{X}$. Furthermore, a sufficiently small neighborhood of any singular point in $U$ is isomorphic to the product of a Kleinian singularity and an open disk. Therefore, by Corollary 2.14, the actions $\lambda_{X}$ and $j_{X}$ lift to the preimage $\pi^{-1}(U)\subset Y$. Since the crepant resolution $\pi$ is semi-small, the complement $Y\setminus\pi^{-1}(U)$ has codimension at least $2$ in $Y$. Thus, the claim follows from Hartogs’ extension theorem. ∎
@@ 2.3. Universal Poisson deformations
In this section, we show that when a conical symplectic variety with a good triple admits a crepant resolution, the good triple naturally extends to the universal Poisson deformation of the resolution.
@@
Theorem 2.15 (Namikawa [14, §5], [13, 15] ).
Let $X$ be a conical symplectic variety with a good triple $(\lambda,\omega,j)$. Assume that $X$ admits a crepant resolution $Y$. Then, there exist universal Poisson deformations $\mathcal{X}\rightarrow\mathcal{B}$ and $\mathcal{Y}\rightarrow\mathcal{C}$ of $X$ and $Y$, respectively. Here, the base spaces $\mathcal{B}$ and $\mathcal{C}$ are vector spaces, and $\mathcal{C}$ is naturally isomorphic to $H^{2}(Y;\mathbb{C})$ via the period map. Furthermore, there exists a quotient map $q:\mathcal{C}\rightarrow\mathcal{B}$ by a finite group such that if we let $\mathcal{X}^{\prime}\rightarrow\mathcal{C}$ be the pullback of $\mathcal{X}$ via $q$, then $\mathcal{Y}$ is a simultaneous resolution of $\mathcal{X}^{\prime}$. That is, there exists a morphism $\nu:\mathcal{Y}\rightarrow\mathcal{X}^{\prime}$ such that the following diagram commutes:
@@
Theorem 2.17 (Namikawa [16], [13, Prop.A.7]).
Let $(X,\lambda_{X},\omega_{X})$ be a conical symplectic variety. Assume that $X$ admits a crepant resolution $Y$. Let $\mathcal{X}\rightarrow\mathcal{B}$ and $\mathcal{Y}\rightarrow\mathcal{C}$ be the universal Poisson deformations of $X$ and $Y$, respectively. Then, the $\mathbb{C}^{*}$-action $\lambda_{X}$ on $X$ uniquely extends to a $\mathbb{C}^{*}$-action $\tilde{\lambda}_{X}$ on the deformation $\mathcal{X}$. Moreover, this $\mathbb{C}^{*}$-action $\tilde{\lambda}_{X}$ naturally lifts to a $\mathbb{C}^{*}$-action $\tilde{\lambda}_{Y}$ on the deformation $\mathcal{Y}$ making the diagram (2.1) commutative. Furthermore, the $\mathbb{C}^{*}$-action induced by $\tilde{\lambda}_{Y}$ on the base space $\mathcal{C}$ is scalar multiplication of weight $2=\mathrm{wt}(\omega)$.
@@
Let $X$ be a conical symplectic variety with a good triple. Assume that $X$ admits a crepant resolution $Y$. Let $\mathcal{Y}\rightarrow\mathcal{C}$ be the universal Poisson deformation of the crepant resolution $Y$. Then, the affinization $\mathcal{X}^{\prime}\rightarrow\mathcal{C}$ of $\mathcal{Y}$ admits a good triple naturally induced from the good triple on $X$. Moreover, this good triple naturally lifts to the triple on $\mathcal{Y}$.
@@
we prove the following lemma on the uniqueness of morphisms on formal Poisson deformations:
@@
Let $X$ be a conical symplectic variety. Let $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ be two Poisson morphisms on a formal Poisson deformation $\{(X_{n}\rightarrow S_{n},\omega_{n})\}_{n}$ of $X$. Assume they satisfy the following three conditions:
@@
    They agree on the central fiber $X$. That is, $f_{n}(0)=f_{n}^{\prime}(0)$.
@@
    They agree on the base space $S_{n}$. That is, for the projection $\pi_{n}:X_{n}\rightarrow S_{n}$, we have $\pi_{n}\circ f_{n}=\pi_{n}\circ f_{n}^{\prime}$.
@@
Since $H^{1}(G,A_{n})=0$ by Lemma 2.20, $\{u_{g}\}_{g\in G}$ is a $1$-coboundary. Therefore, there exists an element $\theta_{n}\in A_{n}$ such that, in particular, $u_{\varepsilon}={}^{\varepsilon}\theta_{n}\circ\theta_{n}^{-1}$. Rearranging this, it follows that $f_{n}$ and $f_{n}^{\prime}$ are conjugate. Furthermore, since the Poisson deformation functor of $X$ is pro-representable, one can verify that we can choose $\theta_{n}$ to be an extension of $\theta_{n-1}$ (cf. Proof of Prop. 2.23 (2)). Thus, the two morphisms $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ on the formal Poisson deformation are conjugate. ∎
@@
In the assumption of Lemma 2.21, assume further that the formal Poisson deformation admits a $\mathbb{C}^{*}$-action and that the two morphisms $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ are equivariant with respect to this $\mathbb{C}^{*}$-action. By replacing the vector space $A_{n}$ used in the proof of Lemma 2.21 with its $\mathbb{C}^{*}$-invariant subspace $A_{n}^{\mathbb{C}^{*}}$, a similar argument shows that $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ are $\mathbb{C}^{*}$-equivariantly conjugate.
@@
Here, $(X_{0},\omega_{0})=(X,\omega)$. We prove the claim by induction on the subscript $n$. For n1n-1, assume that a quaternionic structure $j_{n-1}$ on $X_{n-1}$ has been obtained. We show the following four claims in order.
@@
    be the set of Poisson automorphisms of $\eta_{n}$ whose restrictions to $\eta_{n-1}$ are the identity. We denote the Poisson morphism induced by $j_{n}$ as $j_{n}:\eta\rightarrow\bar{\eta}_{n}$. By composing $j_{n}$ with the complex conjugation $\overline{(\cdot)}:\bar{\eta_{n}}\rightarrow\eta_{n}$, we obtain a Poisson morphism $\hat{j}_{n}:\eta_{n}\rightarrow\eta_{n}$. We aim to modify $j_{n}$ so that $\hat{j}_{n}\circ\hat{j}_{n}=-\mathrm{id}_{\eta_{n}}$.
@@
Let $X$ be a conical symplectic variety with a good triple. Then, the good triple on $X$ naturally induces the good triple on the universal Poisson deformation $\mathcal{X}\rightarrow\mathcal{B}$ of $X$.
@@
Restatement of the claim: Let $X$ be a conical symplectic variety with a good triple. Assume that $X$ admits a crepant resolution $Y$. Let $\mathcal{Y}\rightarrow\mathcal{C}$ be the universal Poisson deformation of the crepant resolution $Y$. Then, the affinization $\mathcal{X}^{\prime}\rightarrow\mathcal{C}$ of $\mathcal{Y}$ admits the good triple naturally induced from the good triple on $X$. Moreover, this good triple naturally lifts to the triple on $\mathcal{Y}$.
We give the proof below. By Corollary 2.25, we obtain the good triple $(\lambda,\omega,j)$ on the universal Poisson deformation $\mathcal{X}\rightarrow\mathcal{B}$ of $X$. The deformation $\mathcal{X}^{\prime}\rightarrow\mathcal{C}$ is, by definition, the Poisson deformation obtained as the pullback of $\mathcal{X}\rightarrow\mathcal{B}$ via the quotient map $q:\mathcal{C}\rightarrow\mathcal{B}$ (cf. Theorem 2.15). As seen in the proof (1) of Proposition 2.23, the quaternionic structure $j$ induces a real structure on the base space $\mathcal{B}$. Thus, one can choose a real structure $\sigma_{\mathcal{C}}$ on $\mathcal{C}$ that is compatible with the quotient map $q$. With this choice, using the universality of the pullback of the deformation $\mathcal{X}^{\prime}$, we obtain the good triple $(\lambda^{\prime},\omega^{\prime},j^{\prime})$ on $\mathcal{X}^{\prime}$. Next, we show that the good triple $(\lambda^{\prime},\omega^{\prime},j^{\prime})$ on the deformation $\mathcal{X}^{\prime}$ lifts to its simultaneous resolution $\mathcal{Y}\rightarrow\mathcal{C}$. By Theorem 2.15, the exceptional set of the simultaneous resolution $\nu:\mathcal{Y}\rightarrow\mathcal{X}^{\prime}$ has codimension at least 2. Therefore, by Hartogs’ extension theorem, it follows that the triple lifts to the deformation $\mathcal{Y}\rightarrow\mathcal{C}$. ∎

### 12 Mar 2026 v2

@@ 1. Introduction
A hyperkähler metric is defined as a metric that is Kähler with respect to three complex structures satisfying the quaternionic relations. In this case, we naturally obtain a deformation of holomorphic symplectic manifolds indexed by the Riemann sphere $\mathbb{P}^{1}$, endowed with a real structure and a family of rational curves called twistor lines; this collection of data is referred to as the twistor space. Conversely, when a twistor space is given, the hyperkähler metric can be reconstructed from its structure [5]. From this viewpoint, Hitchin [4] constructed ALE gravitational instantons, which are typical examples of algebraic hyperkähler metrics with asymptotic behavior.
@@
We define a twistor model to be a structure satisfying all conditions of a twistor space except for the existence of twistor lines. As a generalization of twistor models, we define a new notion called the principal twistor model, which consists of a deformation of holomorphic symplectic manifolds indexed by a vector bundle $\mathcal{C}(2)$ over $\mathbb{P}^{1}$ and a real structure on it. For any choice of a real section $s$ of the vector bundle $\mathcal{C}(2)$ (cf. Def. 3.8), we obtain a twistor model $Z_{s}$ by slicing the principal twistor model along the section $s$ (cf. Prop. 3.9). In other words, the twistor model $Z_{s}$ is given by restricting the principal twistor model to the real section $s$. As will be explained later, the principal twistor model provides a natural framework for studying the moduli space of algebraic hyperkähler structures asymptotic to an algebraic hyperkähler cone (Cor. 1.5). In this framework, the asymptotic behavior of the metric can be interpreted as a degeneration of twistor spaces (cf. Remark 4.11).
In this paper, as the asymptotic hyperkähler cone, we consider a conical symplectic variety $X$ with a good triple (cf. Def. 2.5) —that is, an affine variety equipped with a good $\mathbb{C}^{*}$-action, a holomorphic symplectic form, and a quaternionic structure. As will be explained later, this condition is necessary for a conical symplectic variety to admit an algebraic hyperkähler cone metric (Theorem 1.3). We further assume that the variety $X$ admits a projective crepant resolution $Y$ (which we simply call a crepant resolution).
To study the geometry of $Y$, we use the theory of Poisson deformations. A Poisson deformation is a deformation of a complex manifold equipped with a compatible deformation of the Poisson bracket, which corresponds to a relative holomorphic symplectic form. By a theorem of Namikawa [11], the crepant resolution $Y$ admits the universal Poisson deformation $\mathcal{Y}\to\mathcal{C}$, which captures all such deformations, and its base space $\mathcal{C}$ is naturally isomorphic to $H^{2}(Y;\mathbb{C})$. Based on the idea of Bielawski and Foscolo [2] of gluing this universal Poisson deformation via the $\mathbb{C}^{*}$-action, the principal twistor model $\mathcal{Y}(1)$ is constructed (Prop. 1.1; see Figures 2 and 2 on page 2).
We present three main results of this paper. First, we state that the principal twistor model is naturally constructed from a good triple on a conical symplectic variety. This result is shown by arguments on formal universal Poisson deformations and applying a $\mathbb{C}^{*}$-gluing construction (cf. §3.2):
@@
Let $X$ be a conical symplectic variety with a good triple, and assume that it admits a crepant resolution $Y$. Let $\mathcal{Y}\rightarrow\mathcal{C}$ denote the universal Poisson deformation of $Y$, where the base space $\mathcal{C}$ is isomorphic to $H^{2}(Y;\mathbb{C})$. Then, the good triple naturally extends to a triple on $\mathcal{Y}$. Furthermore, by applying the $\mathbb{C}^{*}$-gluing construction to this triple, the principal twistor model $\mathcal{Y}(1)\rightarrow\mathcal{C}(2)$ is naturally constructed.
@@
The extension of the data other than the quaternionic structure to the universal Poisson deformation is due to Namikawa [10, 13]. In this work, we show that the quaternionic structure can be similarly extended.
The following theorem on the universality of the principal twistor model is the central result of this paper. This theorem states that any twistor space corresponding to an algebraic hyperkähler metric asymptotic to a hyperkähler cone is uniquely obtained by slicing the principal twistor model along a real section. This result is shown by arguments on the universality of the universal Poisson deformation and the asymptotic behavior of the metric:
@@
Let $X$ be a conical symplectic variety, and assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Then, the metric $g_{0}$ naturally defines a good triple on $X$. Suppose further that $X$ admits a crepant resolution $Y$. Let $\mathcal{Y}(1)\rightarrow\mathcal{C}(2)$ be the principal twistor model constructed in Proposition 1.1 for this good triple. Consider an algebraic hyperkähler metric $g$ on $Y$ (cf. Def. 3.29) that is asymptotic at infinity to the metric $g_{0}$ (cf. Def. 3.37). Then, there exists a unique real section $s$ of the vector bundle $\mathcal{C}(2)$ such that the twistor space $Z$ corresponding to the metric $g$ is isomorphic, as a twistor model, to the model $Z_{s}$ obtained by slicing the principal twistor model $\mathcal{Y}(1)$ along the real section $s$.
@@
Finally, as a corollary of Theorem 1.3, we present an application to the moduli space of hyperkähler structures with asymptotic behavior. This result is obtained by combining an analysis of twistor lines with the asymptotic behavior of the metric:
@@
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and that $X$ has a crepant resolution $Y$. Let $\mathcal{M}$ be the moduli space of hyperkähler structures $(Y,g,I,J,K)$ defined as follows:
@@
The organization of the paper is as follows. In §2, we introduce conical symplectic varieties with a good triple and show that the triple extends to the universal Poisson deformation of its crepant resolution. In §3, we define the principal twistor model and explain its construction after introducing the $\mathbb{C}^{*}$-gluing construction. We also prove the universality of the principal twistor model. In §4, as an application of this work, we study the moduli space of hyperkähler structures with asymptotic behavior. We also introduce metrics constructed via hyperkähler quotients and QALE hyperkähler metrics as concrete examples to which this work applies under certain assumptions. Finally, we conclude the paper by discussing future directions.
@@ 2.1. Definitions
Next, we define a quaternionic structure on a complex variety equipped with a $\mathbb{C}^{*}$-action.
@@
An anti-holomorphic automorphism $j$ on a complex variety $X$ with a $\mathbb{C}^{*}$-action $\lambda$ is called a quaternionic structure on $X$ if it satisfies
@@
    The vector space $\mathbb{C}^{2n}$ admits a standard good triple as follows: in terms of standard coordinates $(x_{1},\ldots,x_{n};y_{1},\ldots,y_{n})$,
@@
    Therefore, $X$ naturally inherits the good triple from the standard one on $\mathbb{C}^{2n}$.
@@ 2.2. Universal Poisson deformations
In this subsection, we show that when a conical symplectic variety with a good triple admits a crepant resolution, the good triple naturally extends to the universal Poisson deformation of the resolution.
@@
Theorem 2.10 (Namikawa [11, §5], [10, 12] ).
Let $X$ be a conical symplectic variety. Assume that $X$ admits a crepant resolution $Y$. Then, there exist the universal Poisson deformations $\mathcal{X}\rightarrow\mathcal{B}$ and $\mathcal{Y}\rightarrow\mathcal{C}$ of $X$ and $Y$, respectively. Here, the base spaces $\mathcal{B}$ and $\mathcal{C}$ are vector spaces, and $\mathcal{C}$ is naturally isomorphic to $H^{2}(Y;\mathbb{C})$ via the period map. Furthermore, there exists a quotient map $q:\mathcal{C}\rightarrow\mathcal{B}$ by a finite group such that if we let $\mathcal{X}^{\prime}\rightarrow\mathcal{C}$ be the pullback of $\mathcal{X}$ via $q$, then $\mathcal{Y}$ is a simultaneous resolution of $\mathcal{X}^{\prime}$. That is, there exists a morphism $\nu:\mathcal{Y}\rightarrow\mathcal{X}^{\prime}$ such that the following diagram commutes:
@@
Theorem 2.12 (Namikawa [13], [10, Prop.A.7]).
Let $(X,\omega)$ be a conical symplectic variety. Assume that $X$ admits a crepant resolution $Y$. Let $\mathcal{X}\rightarrow\mathcal{B}$ and $\mathcal{Y}\rightarrow\mathcal{C}$ be the universal Poisson deformations of $X$ and $Y$, respectively. Then, the $\mathbb{C}^{*}$-action $\lambda$ on $X$ uniquely extends to a $\mathbb{C}^{*}$-action $\tilde{\lambda}$ on the deformation $\mathcal{X}$. Moreover, this $\mathbb{C}^{*}$-action $\tilde{\lambda}$ naturally lifts to a $\mathbb{C}^{*}$-action $\tilde{\lambda}_{Y}$ on the deformation $\mathcal{Y}$ making the diagram (2.1) commutative. Furthermore, the $\mathbb{C}^{*}$-action induced by $\tilde{\lambda}_{Y}$ on the base space $\mathcal{C}$ is scalar multiplication of weight $2=\mathrm{wt}(\omega)$.
@@
[…]
@@
We prove the following lemma on the uniqueness of automorphisms on formal Poisson deformations:
@@
Let $X$ be a conical symplectic variety. Let $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ be two Poisson automorphisms on a formal Poisson deformation $\{(X_{n}\rightarrow S_{n},\omega_{n})\}_{n}$ of $X$. Assume they satisfy the following three conditions:
@@
    They agree on the central fiber $X$. That is, $f_{0}=f^{\prime}_{0}$.
@@
    They agree on the base space $S_{n}$ for each $n$. That is, for the projection $\pi_{n}:X_{n}\rightarrow S_{n}$, we have $\pi_{n}\circ f_{n}=\pi_{n}\circ f_{n}^{\prime}$.
@@
Since $H^{1}(G,A_{n})=0$ by Lemma 2.15, $\{u_{g}\}_{g\in G}$ is a $1$-coboundary. Therefore, there exists an element $\theta_{n}\in A_{n}$ such that, in particular, $u_{\varepsilon}={}^{\varepsilon}\theta_{n}\circ\theta_{n}^{-1}$. Rearranging this, it follows that $f_{n}$ and $f_{n}^{\prime}$ are conjugate. Furthermore, since the Poisson deformation functor of $X$ is pro-representable, one can verify that we can choose $\theta_{n}$ to be an extension of $\theta_{n-1}$ (cf. Proof of Prop. 2.18 (2)). Thus, the two morphisms $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ on the formal Poisson deformation are conjugate. ∎
@@
In the assumption of Lemma 2.16, assume further that the formal Poisson deformation admits a $\mathbb{C}^{*}$-action and that the two automorphisms $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ are equivariant with respect to this $\mathbb{C}^{*}$-action. By replacing the vector space $A_{n}$ used in the proof of Lemma 2.16 with its $\mathbb{C}^{*}$-invariant subspace $A_{n}^{\mathbb{C}^{*}}$, a similar argument shows that $\{f_{n}\}_{n}$ and $\{f_{n}^{\prime}\}_{n}$ are $\mathbb{C}^{*}$-equivariantly conjugate.
@@
where $(X_{0},\omega_{0})=(X,\omega)$. We prove the claim by induction on the subscript $n$. For n1n-1, assume that a quaternionic structure $j_{n-1}$ on $X_{n-1}$ has been obtained. We show the following four claims in order.
@@
    be the set of Poisson automorphisms of $\eta_{n}$ whose restrictions to $\eta_{n-1}$ are the identity. We denote the Poisson morphism induced by $j_{n}$ as $j_{n}:\eta_{n}\rightarrow\bar{\eta}_{n}$. By composing $j_{n}$ with the complex conjugation $\overline{(\cdot)}:\bar{\eta_{n}}\rightarrow\eta_{n}$, we obtain a Poisson morphism $\hat{j}_{n}:\eta_{n}\rightarrow\eta_{n}$. We aim to modify $j_{n}$ so that $\hat{j}_{n}\circ\hat{j}_{n}=-\mathrm{id}_{\eta_{n}}$.
@@
[…]
@@
[…]
@@
Remark 2.22.
By restricting the triple $(\tilde{\lambda}_{Y},\tilde{\omega}_{Y},\tilde{j}_{Y})$ on $\mathcal{Y}$ to the central fiber $Y$, we obtain a triple $(\lambda_{Y},\omega_{Y},j_{Y})$. Consequently, this implies that the good triple on $X$ naturally lifts to a triple on the crepant resolution $Y$.